The width w of a rectangular swimming pool is 7x^2. The area A of the pool is 7x^3 – 42x^2. What is an expression for the length of the pool?
Answer by nerdybill(7384)

The width w of a rectangular swimming pool is 7x^2. The area A of the pool is 7x^3 – 42x^2. What is an expression for the length of the pool?
.
width(w) = 7x^2
Area(A) = 7x^3 – 42x^2 = 7x^2(x-6)
.
Since:
Area = width*length
Area/width = length
or
[7x^2(x-6)]/[7x^2] = length
(x-6) = length
x-6 = length (this is what they're looking for)